OpenStudy (anonymous):

What is the solution to the system of equations? 2x + y = -3 x - 2y= -4

OpenStudy (tkhunny):

Have you considered solving it? Personally, I would multiply the first equation by 2 and see if anything magical occurs.

OpenStudy (anonymous):

i just dont get the math

OpenStudy (tkhunny):

That is false. You get lots of math. Now, multiply everything in that first equation by 2 and we can talk.

OpenStudy (anonymous):

We can start by eliminating y, or at list representing it with it's relative value to x. Just shift around the equation 2x+y=-3 y=3-2x There, now we have y And we can just put it into the second equation. x-2(3-2x)=-4 Now just solve this one. :)

OpenStudy (anonymous):

if x-2(3-2x)=-4 is a little confusing, you can simplify it down to 5x-6=-4 This can easily be re-arranged to find x.